OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x2 - 4x + 6. answer choices are ... (-2, 18) (2, 2) (2, 6)

OpenStudy (solomonzelman):

I will re-write in the equation for you in a standard form. \(\normalsize\color{blue}{ y = x^2 - 4x + 6}\) \(\normalsize\color{blue}{ y = x^2 - 4x + 4+2 }\) \(\normalsize\color{blue}{ y = (x-2)^2+2 }\)

OpenStudy (solomonzelman):

I want some effort on your behalf, please.

OpenStudy (solomonzelman):

I would just plug in the options, and only 1 choice would work.